OpenStudy (anonymous):

sara's mother is three times as old as Sara, and 14 years from now she will be twice as old as Sara will be then. how old is each now? i need help figuring out how to write the equation!!!!!!

OpenStudy (anonymous):

@unklerhaukus

mathslover (mathslover):

Let the age of sara = x Age of sara's mother = 3 * age of sara = 3 * x = 3x 14 years from now, Age of sara's mother will be : 3x + 14 Sara's age will be : x + 14 Given, after 14 years , sara's mother's age = 2 * sara's age (3x+14) = 2(x+14) 3x+14 = 2x+28 Solve it now.

OpenStudy (jack1):

much better, my bad

OpenStudy (anonymous):

so it will be 1x+14=28?

mathslover (mathslover):

Yes, solve for x from there.

OpenStudy (anonymous):

so x=14

mathslover (mathslover):

Yep! Right. Good work

mathslover (mathslover):

Also, Welcome to OpenStudy brenroze.

OpenStudy (anonymous):

thanks so much

mathslover (mathslover):

You're welcome. Keep asking questions and do not forget to help others also.

OpenStudy (anonymous):

alright